# The relevance of ‘on’ and ‘off’ transitions in quantum pair production experiments

### v1

@@ V Schwinger effect
For a homogeneous, time-dependent electric field $\textbf{E}(t)$, the temporal gauge $A_{\mu}(t,\bm{x})=(0,\textbf{A}(t))$ simplifies the equations of motion, making them homogeneous. In this gauge, the electric field is given by $\textbf{E}(t)=-\dot{\textbf{A}}(t)$. We consider that the direction of the electric field remains fixed over time. Without loss of generality, we align it along the $z$-axis.
@@
Since the frequency $\Omega_{\bm{k}}$ is independent of any time derivatives of the potential, the analysis regarding its time variation yields conclusions similar to those in the case of a cosmologically conformally coupled scalar field. However, the intermediate regime has a more significant impact on the particle spectrum in the Schwinger effect than in the cosmological case. As shown in Fig. 5, the total number of produced particles in the Schwinger effect continues to increase for large values of $t_{\text{off}}$, in contrast to the cosmological case (Fig. 2), where particle production eventually converges. This is a consequence of the anisotropic nature of the Schwinger effect, which causes the intermediate regime to become increasingly dominant over abrupt transitions as the electric field remains switched on for a sufficiently long duration. Indeed, the linear dependence on $k$ in the frequency (20) causes higher $k$ modes to become excited as the duration of the electric field increases, particularly for modes aligned with the electric field direction. This is consistent with Ref. [52], where it is demonstrated that, when the electric field remains on for sufficiently long durations, the dominant contribution to particle production comes from the intermediate regime rather than the switch-on and switch-off transitions. Nonetheless, even in this case—and even more so when the electric field is switched on for shorter durations—the effects of ‘on’ and ‘off’ transitions remain unavoidable.
Figure 5: Total number density of produced particles $n$ as function of $t_{\text{off}}$ ($t_{\text{on}}=0$) in the Schwinger effect for electric potentials of the form (13), with an intermediate electric strength of $-A_{0}^{\prime}$ and fast switch-on and switch-off transitions ($\delta=0.1/A_{0}^{\prime}$).
@@ VI Conclusions
This issue is particularly critical in analog gravity experiments that simulate a non-conformally coupled field to an expanding FLRW universe, such as [53, 25, 27, 28]. We showed that transitions dominate particle production, effectively overshadowing the contributions from the intermediate dynamics. Therefore, one has to be careful when interpreting the outcomes of such experiments, as the main contribution to pair production does not come from the specific expansion during the intermediate phase that the setup is intended to simulate. In scenarios involving alternating periods of expansion and contraction [46, 28, 47], rapid changes in the scale factor significantly affect the particle spectrum, effectively acting as ‘on’ and ‘off’ transitions within the intermediate regime.
@@ Acknowledgements.
The authors would like to thank Mercedes Martín-Benito and Luis J. Garay for helpful discussions. This work is part of the R+D+I Projects PID2022-139841NB-I00 and PID2023-149018NB-C44, funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. Additionally, A.P.L. also acknowledges support through MICIU fellowship FPU20/05603, project PID2021- 127726NB- I00 (MCIU/AEI/FEDER, UE), Grant IFT Centro de Excelencia Severo Ochoa CEX2020- 001007-S, funded by MCIN/AEI/10.13039/501100011033, the CAM/FEDER Project TEC-2024/COM 84 QUITEMAD-CM, and from the CSIC Research Platform on Quantum Technologies PTI-001.

### v2

@@ V Schwinger effect
For a homogeneous, time-dependent electric field $\textbf{E}(t)$, the temporal gauge $A_{\mu}(t,\bm{x})=(0,\textbf{A}(t))$ simplifies the equations of motion, making them homogeneous. In this gauge, the electric field is given by $\textbf{E}(t)=-\dot{\textbf{A}}(t)$. We consider that the direction of the electric field remains fixed over time. Without loss of generality, we align it along the $z$ axis.
@@
Since the frequency $\Omega_{\bm{k}}$ is independent of any time derivatives of the potential, the analysis regarding its time variation yields conclusions similar to those in the case of a cosmologically conformally coupled scalar field. However, the intermediate regime has a more significant impact on the particle spectrum in the Schwinger effect than in the cosmological case. As shown in Fig. 5, the total number density of produced particles in the Schwinger effect continues to increase for large values of $t_{\text{off}}$, in contrast to the cosmological case (Fig. 2), where particle production eventually converges. This is a consequence of the linear dependence on $k$ in the frequency (20) through the anisotropic term $2q|\textbf{A}|\cos\theta\,k$. As the duration of the electric field increases, this term drives the excitation of higher-$k$ modes with $q\cos\theta<0$, especially those aligned with the field direction, which provide the dominant contribution. This causes the intermediate regime to become increasingly dominant over abrupt transitions as the electric field remains switched on for a sufficiently long duration. This is consistent with Ref. [52], where it is demonstrated that, when the electric field remains on for sufficiently long durations, the dominant contribution to particle production comes from the intermediate regime rather than the switch-on and switch-off transitions. Nonetheless, even in this case—and even more so when the electric field is switched on for shorter durations—the effects of ‘on’ and ‘off’ transitions remain unavoidable.
Figure 5: Total number density of produced particles $n$ as function of $t_{\text{off}}$ ($t_{\text{on}}=0$) in the Schwinger effect for electric potentials of the form (13), with an intermediate electric strength equal to the Schwinger limit, $\dot{A}_{0}=m^{2}/q^{2}$, and fast switch-on and switch-off transitions ($\delta=0.1$). Time variables are expressed in units of $m^{-1}$, and the number density in units of $m^{3}$.
@@ VI Conclusions
This issue is particularly critical in analog gravity experiments that simulate a nonconformally coupled field to an expanding FLRW universe, such as [53, 25, 27, 28]. From our analysis of the time dependence of the system’s characteristic frequency, grounded in quantum kinetic arguments, we showed that transitions dominate particle production, effectively overshadowing the contributions from the intermediate dynamics. Therefore, one has to be careful when interpreting the outcomes of such experiments, as the main contribution to pair production does not come from the specific expansion during the intermediate phase that the setup is intended to simulate. In scenarios involving alternating periods of expansion and contraction [46, 28, 47], rapid changes in the scale factor significantly affect the particle spectrum, effectively acting as ‘on’ and ‘off’ transitions within the intermediate regime.
@@ Acknowledgements.
The authors would like to thank Mercedes Martín-Benito and Luis J. Garay for helpful discussions. This work is part of the R+D+I Projects No. PID2022-139841NB-I00 and No. PID2023-149018NB-C44, funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. Additionally, A.P.L. acknowledges support through MICIU fellowship FPU20/05603, project No. PID2021- 127726NB- I00 (MCIU/AEI/FEDER, UE), Grant IFT Centro de Excelencia Severo Ochoa No. CEX2020- 001007-S, funded by MCIN/AEI/10.13039/501100011033, the CAM/FEDER Project No. TEC-2024/COM 84 QUITEMAD-CM, and from the CSIC Research Platform on Quantum Technologies PTI-001.